# MeSS: City Mesh-Guided Outdoor Scene Generation with Cross-View Consistent Diffusion

### v2

@@ 1 Introduction
Projects like 3DCityDB (Yao, Chaturvedi, and Kolbe 2025), Hamburger LOD3 (Landesbetrieb Geoinformation und Vermessung Hamburg 2025), Ingolstadt LOD3 (SAVeNoW 2020), and 3DBAG (Peters et al. 2022) provide detailed city-scale mesh models, typically at Level of Detail 3 (LoD3) for buildings. However, these models often lack high quality textures, relying instead on simple templates—or omitting textures entirely. This limitation hinders many downstream applications. For example, in VR-based urban navigation, texture quality is essential to user experience—especially at street level, where users are close enough to notice visual flaws.
@@
Recent works like (Chai et al. 2023; Lin et al. 2023; Chen, Wang, and Liu 2023; Xie et al. 2024a, b, 2025) have investigated to generate a city as NeRF (Mildenhall et al. 2021) or 3DGS (Kerbl et al. 2023) representations from city semantic & height map. Since their training process is supervised by images in birds-eye-view, their visual quality degrades significantly after zooming-in. This in turn has fueled efforts to synthetically generate visual data from the perspective on the street.
To alleviate the cost and complexity of real-world data collection for autonomous driving, recent approaches (Wang et al. 2024b; Li, Zhang, and Ye 2024; Zhao et al. 2024a; Gao et al. 2025) have explored generating driving scenes directly from video diffusion models (Blattmann et al. 2023). These methods primarily focus on synthesizing foreground elements such as traffic participants and road infrastructure, leaving the background—particularly buildings—to be hallucinated by the generative model without explicit structural constraints. As a result, the generated videos often fail to align with the actual city layout or urban topology.
To tackle this limitation, Streetscapes (Deng et al. 2024) leverages simplified building geometries and map layouts as conditioning inputs to enhance geometric consistency. Similarly, Cosmos-Transfer1 (Alhaija et al. 2025) integrates background structures into the generation process by utilizing sparse depth data derived from lidar point clouds, though its depth control mechanism ignores all scene content beyond a 75-meter range. Nevertheless, due to the imprecise nature of their geometric inputs and the feedforward design of their conditional generation pipelines, these methods fail to guarantee precise geometric alignment.
@@
Directly texturing mesh models with 2D diffusion models  (Chen et al. 2023a; Metzer et al. 2023; Chen et al. 2024; Richardson et al. 2023) may seem a viable option, but existing methods are limited to single objects or small scenes with low polygon counts, far from city-scale street scenes. Another approach is perpetual view generation from a single image (Yu et al. 2024a, b; Wang et al. 2024a; Chung et al. 2023), which relies on progressively outpainting of RGB images and associated depth maps to cover the scene, using 2D diffusion models (Rombach et al. 2022). Their outpainting strategy often gradually drifts away from the initial appearance and is not capable for generating consistent dense views.
@@ 2.1 Scene Generation from 2D Map
InfiniCity (Lin et al. 2023) decomposes 3D scene generation into three steps: 2D map generation, map-to-voxel lifting, and voxel texturing through neural rendering. SceneDreamer (Chen, Wang, and Liu 2023) represents unbounded 3D landscapes via a bird’s-eye view (BEV) layout with height and semantic fields, manipulated through 2D generative models. Recent extensions like CityDreamer (Xie et al. 2024a) and its temporal variant CityDreamer4D (Xie et al. 2025) refine this paradigm by decomposing the rendering process into specialized modules for distinct scene components. CityDreamer4D explicitly separates static backgrounds, dynamic buildings, and movable foreground objects, while incorporating temporal consistency for 4D urban scene synthesis. Despite these advancements, existing methods (Xie et al. 2024b, a, 2025) still face challenges in stabilizing layout generation and achieving high-fidelity rendering.
@@ 2.2 Perpetual View Generation
Infinite-Nature (Liu et al. 2021) defines perpetual view generation as synthesizing views along arbitrary camera trajectories from a single image, employing a render-and-refine pipeline that progressively extends scenes via outpainting with cross-frame depth alignment. This principle has been extended through GANs (Li et al. 2022) and diffusion models: DiffDreamer (Cai et al. 2023) and SceneScape (Fridman et al. 2023) focus on static scene extrapolation, while WonderWorld (Yu et al. 2024a) enables interactive 3D scene generation through guided depth diffusion, allowing users to specify scene contents and viewpoints. For view-consistent generation, VistaDream (Wang et al. 2024a) explicitly enforces multiview geometry constraints during denoising process. These and other similar approaches (Chung et al. 2023; Yu et al. 2024b) all profit from advances in monocular depth estimation (Ranftl et al. 2020; Bhat et al. 2023; Ke et al. 2024a). However, they still exhibit rapid appearance drift in complex scenarios like urban driving, where error accumulation in the outpainting step amplifies inconsistencies across occluded regions.
@@ 2.3 Scene Synthesis from Video Generation
In the context of scene construction, video generators offer a way to synthesize (approximately) consistent views, either by conditioning on a single frame (Yu et al. 2024c; Fan et al. 2024; Sun et al. 2024) or by interpolating multiple given viewpoints (Liu et al. 2024; Yu et al. 2024c; Sun et al. 2024). In particular, StreetScapes (Deng et al. 2024) addresses a task similar to ours with a two-frame video generator based on AnimateDiff (Guo et al. 2023a). Rendered depth, height, and semantics are used as control inputs to generate consecutive street views. Despite starting from a pretrained Latent Diffusion Model (Rombach et al. 2022), training such a video generator is expensive in terms of training data and compute. The insufficient geometric controling ability also leads to adjacent-frame inconsistency and temporal drift in long sequence.
@@ 3.1 Preliminaries
We utilize a Latent Consistency Model (LCM) to enhance the inpainting consistency across views in our method. Its core idea is to learn a function that maps any point on a trajectory of the probability flow ODE (Song et al. 2020; Lu et al. 2022) to that trajectory’s origin (i.e., the solution of the ODE). LCM is trained by enforcing the self-consistency property with a consistency function $\boldsymbol{f_{\theta}}$:
@@
with $\delta$ a fixed, small positive number and $T$ refers total diffusion steps. For ${\boldsymbol{\epsilon}}\text{-prediction}$ (Song et al. 2020), the function ${\boldsymbol{f}}_{\theta}$ is parameterized as
@@ 3.2 Stage I: Gaussian Field Construction from Key Views
Two multi-conditioned ControlNets (Zhang, Rao, and Agrawala 2023) are introduced to generate $N$ key views in autoregressive fashion, moving backwards from the end to the start of the camera path***Views are indexed in driving direction.. First, ControlNet-S takes control signals $\{\boldsymbol{d}^{N},\boldsymbol{s}^{N},\boldsymbol{n}^{N}\}$ (depth, semantics and normals) at camera pose $\mathcal{C}^{N}$ to generate view image $\boldsymbol{I}^{N}$. As single-channel disparity maps may lose the ability to discriminate depth values in the far field, so we encode depth as a colormap. This ensures depth variations are retained both near the camera and in the far field.
@@
To avoid the computational cost associated with Gaussian Field optimization, splats are directly positioned on the known 3D surface by lifting image pixels according to the depth map. Furthermore, following the approach in SUGAR(Guédon and Lepetit 2024), the Gaussians are aligned with the surface and flattened by reducing the scale factor along the surface normal to a small value. The other two scale parameters are selected to ensure full coverage of the mesh surface, thereby preventing the formation of Moiré patterns and gaps (Yu et al. 2024a). For the Gaussians, Spherical Harmonics coefficients are computed from the RGB values of corresponding pixels, while opacity is set based on empirical observations. Additional details are provided in the supplementary material.
@@ 3.3 Stage II: Gaussian Field Densification
We render from above constructed scene at intermediate novel views $\{\mathcal{C}^{l}\}$ between two keyframes $\boldsymbol{I}^{n}$ and $\boldsymbol{I}^{n+1},n\in[1,N)$. As illustrated in Fig. 3, the novel views are inspected with silhouettes since they are rendered from the coarse scene model. Novel views not included during the 3DGS reconstruction commonly suffer from holes and blurry textures (Kerbl et al. 2023). Inspired by (Epstein et al. 2023; Dhariwal and Nichol 2021; Luo et al. 2024; Yu et al. 2024a), we modify the sampling process of latent diffusion to obtain a training-free method for RGB inpainting, termed Appearance Guided Inpainting (AGInpaint).
@@ 3.4 Global Consistency Alignment
Gaussian field generated with the above two stages so far still exhibit appearance artifacts such as brightness drift, outpainting mistakes, etc. To resolve those, we draw inspiration from Multiview Consistency Sampling (MCS) in VistaDream (Wang et al. 2024a) and design a cleaning method called Global Consistency Alignment (GCA), tailored for our sparse-to-dense generation pipeline.
@@ 4.1 Data Preparation and Implementation
To assess the feasibility of our pipeline, we render RGB images with corresponding depth, normal, and semantic maps using City Sample Project (Epic Games 2022) in Unreal Engine 5 (Epic Games 2021). ControlNets are trained upon frozen Stable Diffusion 1.5 (SD1.5) with training data rendered above. More details about data preparation and implementation can be found in the supplementary material.
@@ 4.2 Comparison with other methods
Both WonderWorld  (Yu et al. 2024a) and VistaDream (Wang et al. 2024a) rely on monocular depth estimation methods  (Bochkovskii et al. 2024; Ke et al. 2024b) to give depth value for placing Gaussians into 3D space, while our method is based on mesh geometry with precise metric depth. To form a fair comparison, we inject metric depth into their pipelines and align their inpainting diffusion model with ours. The altered ones are named as WonderWorld and VistaDream respectively. For each camera path, We start the scene from the same initial view and outpaint them by moving the camera backwards for 100 meters using the two and our pipeline. In this way, 3200 images are produced in total.
@@
CityDreamer4D  (Xie et al. 2025) is implemented on the Citytopia dataset  (Xie et al. 2025), which is also constructed based on the City Sample Project in Unreal Engine 5 (UE5), similar to ours. In comparison, our method demonstrates superior performance, as evidenced by lower scores in the FID and KID metrics. Furthermore, as illustrated in Fig. 5, their generated scenes exhibit notable limitations in visual quality upon closer inspection, particularly at street level. Specifically, the results lack critical urban elements such as streetlamps, traffic signs, and other street-level instances, significantly diminishing their realism and practical applicability.
@@
For StreetScapes  (Deng et al. 2024), we directly use the results reported in their paper since their code is not publicly available. However, forming a fair comparison is challenging, as their model is trained on a significantly larger dataset compared to oursEven for their experiment on the London patch, they utilize approximately 20 times more data than we do.. Given that our generated sequence is longer than theirs (200 frames vs. 64 frames), we evaluate their results within the range of 32 to 64 frames. In terms of quality, our results are on par with theirs, while we outperform them in terms of temporal consistency. This advantage can be attributed to our method’s effective utilization of geometric priors, which plays a key role in achieving superior consistency.
@@ 4.3 Stylized rendering
As you can see, the appearance of our generated scene is heavily branded with City Sample style, simply tweaking text prompt during view generation does not work for stylization. This hinders the application of the scene, so we develop a way to customize it through stylized rendering. Given a camera path in the scene, we render video and transfer it to different style via video relightning method TC-Light (Liu et al. 2025) or SDEdit (Meng et al. 2021) with Cosmos-Transfer1 (Alhaija et al. 2025). As depicted in  Fig. 6, both are capable of completing this task by taking the original views as prior, and Cosmos-Transfer1 outperforms TC-Light in visual fidelity. The styled videos can be projected back to the gaussian scene if desired.
@@
Citydreamer4D  (Wang et al. 2024a) - 88.48 0.049
WonderWorld-Geometry  (Wang et al. 2024a) 0.516 75.807 0.076
Vistadream-Geometry  (Yu et al. 2024a) 0.508 72.44 0.073
StreetScapes  (Deng et al. 2024) 0.519 29.93 0.025
@@ 4.4 Ablation Study
For the outpainting task of key views in Stage I, we replace the preceding frame conditioned ControlNet-N by ControlNet-S. To make it different w.r.t. common warp-and-outpaint pipeline  (Yu et al. 2024a, b), AGInpaint is still equipped. As depicted in Tab. 2, the fidelity of generated frames degenerates dramatically. Besides, the consistency is also getting worse due to larger appearance drift accumulated during scene extrapolation.
@@
We replace the AGInpaint component in our method with a simpler approach called Resample  (Deng et al. 2024; Lugmayr et al. 2022), which involves re-adding several rounds of noise at each denoising step to achieve a more homogeneous inpainted result. As demonstrated in Tab. 2, the absence of AGInpaint significantly impacts the generation quality. Through a visual comparison of the inpainted regions produced by AGInpaint and Resample (Fig. 4(a)), we highlight the superior efficacy of AGInpaint. Due to the downsampling of inpainting masks, Resample struggles to fill in slim regions with Latent Diffusion Models, causing noticeable streak patterns.

### v3

@@ 1 Introduction
Projects like 3DCityDB (3dcitydb), Hamburger LOD3 (hamburgLOD3_2025), Ingolstadt LOD3 (ingolstadt3d2020), and 3DBAG (peters2022) provide detailed city-scale mesh models, typically at Level of Detail 3 (LoD3) for buildings. However, these models often lack high quality textures, relying instead on simple templates—or omitting textures entirely. This limitation hinders many downstream applications. For example, in VR-based urban navigation, texture quality is essential to user experience—especially at street level, where users are close enough to notice visual flaws.
@@
Recent works like (chai2023persistent; lin2023infinicity; chen2023scenedreamer; xie2024citydreamer; xie2024gaussiancity; xie2025citydreamer4d) have investigated to generate a city as NeRF (mildenhall2021nerf) or 3DGS (kerbl20233d) representations from city semantic & height map. Since their training process is supervised by images in birds-eye-view, their visual quality degrades significantly after zooming-in. This in turn has fueled efforts to synthetically generate visual data from the perspective on the street.
To alleviate the cost and complexity of real-world data collection for autonomous driving, recent approaches (wang2024drivedreamer; li2024drivingdiffusion; zhao2024drivedreamer; gao2025vista) have explored generating driving scenes directly from video diffusion models (blattmann2023stable). These methods primarily focus on synthesizing foreground elements such as traffic participants and road infrastructure, leaving the background—particularly buildings—to be hallucinated by the generative model without explicit structural constraints. As a result, the generated videos often fail to align with the actual city layout or urban topology.
To tackle this limitation, Streetscapes (deng2024streetscapes) leverages simplified building geometries and map layouts as conditioning inputs to enhance geometric consistency. Similarly, Cosmos-Transfer1 (alhaija2025cosmos) integrates background structures into the generation process by utilizing sparse depth data derived from lidar point clouds, though its depth control mechanism ignores all scene content beyond a 75-meter range. Nevertheless, due to the imprecise nature of their geometric inputs and the feedforward design of their conditional generation pipelines, these methods fail to guarantee precise geometric alignment.
@@
Directly texturing mesh models with 2D diffusion models  (chen2023text2tex; metzer2023latent; chen2024scenetex; richardson2023texture) may seem a viable option, but existing methods are limited to single objects or small scenes with low polygon counts, far from city-scale street scenes. Another approach is perpetual view generation from a single image (yu2024wonderworld; yu2024wonderjourney; wang2024vistadream; chung2023luciddreamer), which relies on progressively outpainting of RGB images and associated depth maps to cover the scene, using 2D diffusion models (Rombach_2022_CVPR). Their outpainting strategy often gradually drifts away from the initial appearance and is not capable for generating consistent dense views.
@@ 2.1 Scene Generation from 2D Map
InfiniCity (lin2023infinicity) decomposes 3D scene generation into three steps: 2D map generation, map-to-voxel lifting, and voxel texturing through neural rendering. SceneDreamer (chen2023scenedreamer) represents unbounded 3D landscapes via a bird’s-eye view (BEV) layout with height and semantic fields, manipulated through 2D generative models. Recent extensions like CityDreamer (xie2024citydreamer) and its temporal variant CityDreamer4D (xie2025citydreamer4d) refine this paradigm by decomposing the rendering process into specialized modules for distinct scene components. CityDreamer4D explicitly separates static backgrounds, dynamic buildings, and movable foreground objects, while incorporating temporal consistency for 4D urban scene synthesis. Despite these advancements, existing methods (xie2024gaussiancity; xie2024citydreamer; xie2025citydreamer4d) still face challenges in stabilizing layout generation and achieving high-fidelity rendering.
@@ 2.2 Perpetual View Generation
Infinite-Nature (liu2021infinite) defines perpetual view generation as synthesizing views along arbitrary camera trajectories from a single image, employing a render-and-refine pipeline that progressively extends scenes via outpainting with cross-frame depth alignment. This principle has been extended through GANs (li2022infinitenature) and diffusion models: DiffDreamer (cai2023diffdreamer) and SceneScape (fridman2023scenescape) focus on static scene extrapolation, while WonderWorld (yu2024wonderworld) enables interactive 3D scene generation through guided depth diffusion, allowing users to specify scene contents and viewpoints. For view-consistent generation, VistaDream (wang2024vistadream) explicitly enforces multiview geometry constraints during denoising process. These and other similar approaches (chung2023luciddreamer; yu2024wonderjourney) all profit from advances in monocular depth estimation (ranftl2020towards; bhat2023zoedepth; ke2024repurposing). However, they still exhibit rapid appearance drift in complex scenarios like urban driving, where error accumulation in the outpainting step amplifies inconsistencies across occluded regions.
@@ 2.3 Scene Synthesis from Video Generation
In the context of scene construction, video generators offer a way to synthesize (approximately) consistent views, either by conditioning on a single frame (yu2024viewcrafter; fan2024instantsplat; sun2024dimensionx) or by interpolating multiple given viewpoints (liu2024reconx; yu2024viewcrafter; sun2024dimensionx). In particular, StreetScapes (deng2024streetscapes) addresses a task similar to ours with a two-frame video generator based on AnimateDiff (guo2023animatediff). Rendered depth, height, and semantics are used as control inputs to generate consecutive street views. Despite starting from a pretrained Latent Diffusion Model (Rombach_2022_CVPR), training such a video generator is expensive in terms of training data and compute. The insufficient geometric controling ability also leads to adjacent-frame inconsistency and temporal drift in long sequence.
@@ 3.1 Preliminaries
We utilize a Latent Consistency Model (LCM) to enhance the inpainting consistency across views in our method. Its core idea is to learn a function that maps any point on a trajectory of the probability flow ODE (song2020score; lu2022dpm) to that trajectory’s origin (i.e., the solution of the ODE). LCM is trained by enforcing the self-consistency property with a consistency function $\boldsymbol{f_{\theta}}$:
@@
with $\delta$ a fixed, small positive number and $T$ refers total diffusion steps. For ${\boldsymbol{\epsilon}}\text{-prediction}$ (song2020score), the function ${\boldsymbol{f}}_{\theta}$ is parameterized as
@@ 3.2 Stage I: Gaussian Field Construction from Key Views
Two multi-conditioned ControlNets (zhang2023adding) are introduced to generate $N$ key views in autoregressive fashion, moving backwards from the end to the start of the camera path***Views are indexed in driving direction.. First, ControlNet-S takes control signals $\{\boldsymbol{d}^{N},\boldsymbol{s}^{N},\boldsymbol{n}^{N}\}$ (depth, semantics and normals) at camera pose $\mathcal{C}^{N}$ to generate view image $\boldsymbol{I}^{N}$. As single-channel disparity maps may lose the ability to discriminate depth values in the far field, so we encode depth as a colormap. This ensures depth variations are retained both near the camera and in the far field.
@@
To avoid the computational cost associated with Gaussian Field optimization, splats are directly positioned on the known 3D surface by lifting image pixels according to the depth map. Furthermore, following the approach in SUGAR(guedon2024sugar), the Gaussians are aligned with the surface and flattened by reducing the scale factor along the surface normal to a small value. The other two scale parameters are selected to ensure full coverage of the mesh surface, thereby preventing the formation of Moiré patterns and gaps (yu2024wonderworld). For the Gaussians, Spherical Harmonics coefficients are computed from the RGB values of corresponding pixels, while opacity is set based on empirical observations. Additional details are provided in the supplementary material.
@@ 3.3 Stage II: Gaussian Field Densification
We render from above constructed scene at intermediate novel views $\{\mathcal{C}^{l}\}$ between two keyframes $\boldsymbol{I}^{n}$ and $\boldsymbol{I}^{n+1},n\in[1,N)$. As illustrated in Fig. 3, the novel views are inspected with silhouettes since they are rendered from the coarse scene model. Novel views not included during the 3DGS reconstruction commonly suffer from holes and blurry textures (kerbl20233d). Inspired by (epstein2023diffusion; dhariwal2021diffusion; luo2024readout; yu2024wonderworld), we modify the sampling process of latent diffusion to obtain a training-free method for RGB inpainting, termed Appearance Guided Inpainting (AGInpaint).
@@ 3.4 Global Consistency Alignment
Gaussian field generated with the above two stages so far still exhibit appearance artifacts such as brightness drift, outpainting mistakes, etc. To resolve those, we draw inspiration from Multiview Consistency Sampling (MCS) in VistaDream (wang2024vistadream) and design a cleaning method called Global Consistency Alignment (GCA), tailored for our sparse-to-dense generation pipeline.
@@ 4.1 Data Preparation and Implementation
To assess the feasibility of our pipeline, we render RGB images with corresponding depth, normal, and semantic maps using City Sample Project (epic2022citysample) in Unreal Engine 5 (epic2021ue5). ControlNets are trained upon frozen Stable Diffusion 1.5 (SD1.5) with training data rendered above. More details about data preparation and implementation can be found in the supplementary material.
@@ 4.2 Comparison with other methods
Both WonderWorld  (yu2024wonderworld) and VistaDream (wang2024vistadream) rely on monocular depth estimation methods  (bochkovskii2024depth; ke2023repurposing) to give depth value for placing Gaussians into 3D space, while our method is based on mesh geometry with precise metric depth. To form a fair comparison, we inject metric depth into their pipelines and align their inpainting diffusion model with ours. The altered ones are named as WonderWorld and VistaDream respectively. For each camera path, We start the scene from the same initial view and outpaint them by moving the camera backwards for 100 meters using the two and our pipeline. In this way, 3200 images are produced in total.
@@
CityDreamer4D  (xie2025citydreamer4d) is implemented on the Citytopia dataset  (xie2025citydreamer4d), which is also constructed based on the City Sample Project in Unreal Engine 5 (UE5), similar to ours. In comparison, our method demonstrates superior performance, as evidenced by lower scores in the FID and KID metrics. Furthermore, as illustrated in Fig. 5, their generated scenes exhibit notable limitations in visual quality upon closer inspection, particularly at street level. Specifically, the results lack critical urban elements such as streetlamps, traffic signs, and other street-level instances, significantly diminishing their realism and practical applicability.
@@
For StreetScapes  (deng2024streetscapes), we directly use the results reported in their paper since their code is not publicly available. However, forming a fair comparison is challenging, as their model is trained on a significantly larger dataset compared to oursEven for their experiment on the London patch, they utilize approximately 20 times more data than we do.. Given that our generated sequence is longer than theirs (200 frames vs. 64 frames), we evaluate their results within the range of 32 to 64 frames. In terms of quality, our results are on par with theirs, while we outperform them in terms of temporal consistency. This advantage can be attributed to our method’s effective utilization of geometric priors, which plays a key role in achieving superior consistency.
@@ 4.3 Stylized rendering
As you can see, the appearance of our generated scene is heavily branded with City Sample style, simply tweaking text prompt during view generation does not work for stylization. This hinders the application of the scene, so we develop a way to customize it through stylized rendering. Given a camera path in the scene, we render video and transfer it to different style via video relightning method TC-Light (liu2025tc) or SDEdit (meng2021sdedit) with Cosmos-Transfer1 (alhaija2025cosmos). As depicted in  Fig. 6, both are capable of completing this task by taking the original views as prior, and Cosmos-Transfer1 outperforms TC-Light in visual fidelity. The styled videos can be projected back to the gaussian scene if desired.
@@
Citydreamer4D  (wang2024vistadream) - 88.48 0.049
WonderWorld-Geometry  (wang2024vistadream) 0.516 75.807 0.076
Vistadream-Geometry  (yu2024wonderworld) 0.508 72.44 0.073
StreetScapes  (deng2024streetscapes) 0.519 29.93 0.025
@@ 4.4 Ablation Study
For the outpainting task of key views in Stage I, we replace the preceding frame conditioned ControlNet-N by ControlNet-S. To make it different w.r.t. common warp-and-outpaint pipeline  (yu2024wonderworld; yu2024wonderjourney), AGInpaint is still equipped. As depicted in Tab. 2, the fidelity of generated frames degenerates dramatically. Besides, the consistency is also getting worse due to larger appearance drift accumulated during scene extrapolation.
@@
We replace the AGInpaint component in our method with a simpler approach called Resample  (deng2024streetscapes; lugmayr2022repaint), which involves re-adding several rounds of noise at each denoising step to achieve a more homogeneous inpainted result. As demonstrated in Tab. 2, the absence of AGInpaint significantly impacts the generation quality. Through a visual comparison of the inpainted regions produced by AGInpaint and Resample (Fig. 4(a)), we highlight the superior efficacy of AGInpaint. Due to the downsampling of inpainting masks, Resample struggles to fill in slim regions with Latent Diffusion Models, causing noticeable streak patterns.